# Quantum Steenrod operations of symplectic resolutions

### v2

@@ 1.2. Main results
Fix a cohomology class $b\in H^{*}(X;\mathbb{F}_{p})$. The quantum Steenrod operations [Fuk97] [Wil20] are defined from $\mathbb{Z}/p$-equivariant counts of (perturbed) holomorphic curves $u:\mathbb{P}^{1}\to X$ from $\mathbb{P}^{1}$ with $p$ marked points along the roots of unity, all incident to (Poincaré dual of) $b$. Following [Lee23], we generalize the quantum Steenrod operations to a setting where the target symplectic manifold $(X,\omega)$ admits Hamiltonian torus actions by $T=(S^{1})^{r}$. For a fixed (equivariant) cohomology class $b\in H^{*}_{T}(X;\mathbb{F}_{p})$ we define a $T$-equivariant quantum Steenrod operation (2.17) which is an endomorphism
of the underlying module of the equivariant quantum cohomology; here $\Lambda$ is the Novikov ring and $t,\theta$ represent the $\mathbb{Z}/p$-equivariant parameters $H^{*}_{\mathbb{Z}/p}(\mathrm{pt};\mathbb{F}_{p})\cong\mathbb{F}_{p}{[\![t,\theta]\!]}$ of degrees $|t|=2,|\theta|=1$.
In the presence of a torus $T$ acting on $X$, there is another set of operations on equivariant quantum cohomology known as the shift operators (2.27) for each choice of a cocharacter $\beta:S^{1}\to T$,
(1.2) $\mathbb{S}_{\beta}:H^{*}_{{T}}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{{T}}(X;\Lambda){[\![t,\theta]\!]}.$
The key property of the shift operator is that it does not give linear map of $H^{*}_{T}(\mathrm{pt}){[\![t,\theta]\!]}$-modules, but rather satisfies a “twisted-linear” condition (2.56). Following [Iri17] and [LJ20], we define the shift operator as a composition of a linear map $S_{\beta}$ (the “bare” shift operators) and a reparametrization map accounting for the failure of linearity (and hence the shift), see 2.27.
Our first main result is the following new compatibility relation of the (equivariant) quantum Steenrod operations $\Sigma_{b}^{T}$ with the shift operators $\mathbb{S}_{\beta}$.
@@
The quantum Steenrod operations $\Sigma_{b}^{T}$ and the shift operators $\mathbb{S}_{\beta}$ commute. Equivalently, the bare shift operators $S_{\beta}$ commute with $\Sigma_{b}^{T}$ up to a shift in the equivariant parameters:
(1.3) $\Sigma_{b}^{T}|_{\lambda\mapsto\lambda-\beta t}\circ S_{\beta}=S_{\beta}\circ\Sigma_{b}^{T}.$
@@
(1.4) $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}:H^{*}_{T}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{T}(X;\Lambda){[\![t,\theta]\!]}.$
@@
For almost all primes $p$, the $p$-curvature $F_{b}$ and the quantum Steenrod operation $\Sigma_{b}^{\mathbf{T}}$ for $b\in H^{2}(X;\mathbb{F}_{p})$ agree.
@@
In particular, the same result with the same proof holds for e.g. generalized Springer resolutions $T^{*}(G/P)$ for parabolic $P\leq G$ and Nakajima quiver varieties of type $A$; in general it would hold under the assumptions that (i) $X^{T}$ is discrete and (ii) the quantum multiplication $b\ \ast$ has simple spectrum. ∎
@@ 1.3. Methods
(1.5) $\Sigma_{b}^{T}|_{\lambda\mapsto\lambda-\beta t}\circ S_{\beta}=S_{\beta}\circ\Sigma_{b}^{T}$
@@
From 1.1, we deduce 1.2. The key observations are the following (2.18, 4.5) that
@@
To show that such an expansion of $F_{b}$ and $\Sigma_{b}$ as polynomials of $h$ is well-defined, we work in a basis of equivariant cohomology provided by the theory of cohomological stable envelopes of [MO19]. For the case of $X=T^{*}(G/B)$, the stable envelopes are given by certain linear combinations of the classes represented by the conormals of Schubert varieties in $G/B$, see [SZ20]. The main technical input is to show the compactness and transversality of the moduli spaces used to define the quantum Steenrod operations and shift operators, when the input and output incidence constraints are given by the stable envelope classes.
@@ 1.4. Related work
The main result of [SW22] was covariant constancy, which establishes the compatibility of the quantum Steenrod operation $\Sigma_{b}$ with the quantum differential connection. In light of the above discussion, our main result 1.1 can be understood as a companion theorem to [SW22] which establishes the compatibility of $\Sigma_{b}$ with the difference connection. There are very few operations which are compatible with both the quantum differential connection and the difference connection from the shift operators, and this constrains the quantum Steenrod operations enough to be equivalent to $p$-curvature.
The $p$-curvature is a fundamental invariant of a connection in positive characteristic, and is an object of classical interest, notably featuring in the Grothendieck–Katz $p$-curvature conjecture [Kat72]. In the recent important work of Etingof–Varchenko [EV24b, EV24a], they develop the theory of periodic pencils of flat connections and prove systematic results about the spectrum of $p$-curvature endomorphisms. Our equivariant quantum connections of conical symplectic resolutions are examples of periodic pencils of flat connections, where periodicity arises from the compatibility with the shift operators. For the main result of this paper, the precise results from the the theory form [EV24b, EV24a] can actually be bypassed, but we expect their theory to be a key ingredient for the further study of equivariant quantum connections in positive characteristic.
@@ 1.5. Organization of the paper
In Section 4 we review the form and properties of the equivariant quantum connections of symplectic resolutions and their $p$-curvature. We also briefly introduce the theory of periodic pencils of flat connections of Etingof–Varchenko [EV24b], in particular their structural result about the eigenvalues of the $p$-curvature for such connections.
@@ 1.6. Acknowledgements
We would like thank Paul Seidel for his continued encouragement. We thank Shaoyun Bai, Zihong Chen, Hunter Dinkins, Pavel Etingof, Vasily Krylov, Davesh Maulik, Dan Pomerleano, Ben Webster, Alexander Varchenko, and Nicholas Wilkins for stimulating discussions at various stages. We especially thank Pavel Etingof for explaining his ongoing work with Alexander Varchenko and suggesting its crucial relevance with the current project, leading to the main result. This work was partially supported by MIT Landis fellowship, by the National Science Foundation through grant DMS-1904997, and by the Simons Foundation through grants 6552299 (the Simons Collaboration on Homological Mirror Symmetry) and 256290 (Simons Investigator).
@@ 2.1.3. Equivariant cohomology
again as the model of $E{\mathbb{Z}/p}$, where $\eta\in{\mathbb{Z}/p}$ acts freely by multiplication of $\zeta=e^{2\pi i/p}$. Denote this action of ${\mathbb{Z}/p}$ by $\sigma$. The space $E{\mathbb{Z}/p}=S^{\infty}$ can be equipped with a ${\mathbb{Z}/p}$-CW complex structure with exactly $p$ many $i$-cells $\Delta^{i},\sigma\Delta^{i},\dots,\sigma^{p-1}\Delta^{i}$ in each dimension $i\geq 0$. The $i$-cell $\Delta^{i}$ has a natural compactification whose top strata in the boundary are given by the union $\bigsqcup_{j=0}^{p-1}\sigma^{j}\Delta^{i-1}$ of $(i-1)$-cells. The images of $\Delta^{i}$ under the quotient $E{\mathbb{Z}/p}\to B{\mathbb{Z}/p}$ induces a CW-structure on $B{\mathbb{Z}/p}$ with unique $i$-cell for every $i\geq 0$. For $\mathbb{F}_{p}$-coefficients, the images of (the compactification of) $\Delta^{i}$ in $B{\mathbb{Z}/p}$ become cycles, additively generating $H_{*}^{\mathbb{Z}/p}(\mathrm{pt};\mathbb{F}_{p})=H_{*}(B{\mathbb{Z}/p};\mathbb{F}_{p})$.
@@
Then $t,\theta$ generate $H^{*}_{{\mathbb{Z}/p}}(\mathrm{pt},\mathbb{F}_{p})\cong\mathbb{F}_{p}[\![t,\theta]\!]$. Additively, the ${\mathbb{Z}/p}$-equivariant cohomology of a point has one generator in each degree $i\geq 0$, which we will denote by
@@
We again take $S^{\infty}$ with the multiplication action of $S^{1}\leq\mathbb{C}^{*}$ as a model of $ES^{1}$, so that $BS^{1}=\mathbb{CP}^{\infty}$ has one $i$-cell for each $i\in 2\mathbb{Z}_{\geq 0}$. Correspondingly, $H^{*}_{S^{1}}(\mathrm{pt};\mathbb{F}_{p})=\mathbb{F}_{p}[\![\lambda]\!]$ for a choice of the generator $\lambda\in H^{2}(\mathbb{CP}^{\infty};\mathbb{F}_{p})$, given by the first Chern class of the universal line bundle $\mathcal{O}_{\mathbb{CP}^{\infty}}(1)$. For a higher-dimensional torus $T=(S^{1})^{r}$, take the product $ET=(S^{\infty})^{k}$ with the product action so that $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})=\mathbb{F}_{p}[\![\lambda_{1},\dots,\lambda_{r}]\!]:=\mathbb{F}_{p}[\![{\lambda}]\!]$.
@@
Note that locally finite cycles in $X$ represent cohomology classes in $H^{*}(X;\mathbb{F}_{p})$ by Poincaré duality, and $T$-invariant locally finite cycles correspondingly represent cohomology classes in $H^{*}_{T}(X;\mathbb{F}_{p})$. For the running example of $X=T^{*}(G/B)$, it is classical that its cohomology over $\mathbb{Z}$ is additively generated by locally finite cycles given by the unions of conormals to Schubert subvarieties of $G/B$ [Bri05], see Section 3.
@@ 2.1.4. Equivariant quantum cohomology
For closed $X$, the ordinary cohomology of $X$ admits the nondegenerate Poincaré pairing $(\cdot,\cdot):H^{*}(X;\mathbb{F}_{p})\otimes H^{*}(X;\mathbb{F}_{p})\to\mathbb{F}_{p}$ given by $(b,b^{\prime})=\int_{X}b\smile b^{\prime}$. For a fixed $b\in H^{*}(X;\mathbb{F}_{p})$, we denote its linear dual under the Poincaré pairing by $b^{\vee}\in H^{*}(X;\mathbb{F}_{p})$. If $X$ admits an action of $T$, the Poincaré pairing can be extended to the equivariant cohomology ring $(\cdot,\cdot)_{T}:H^{*}_{T}(X)\otimes H^{*}_{T}(X)\to H^{*}_{T}(\mathrm{pt})$.
If $X$ is non-compact, the Poincaré pairing becomes a pairing between cohomology and compactly supported cohomology, $(\cdot,\cdot):H^{*}(X;\mathbb{F}_{p})\otimes H^{*}_{c}(X;\mathbb{F}_{p})\to\mathbb{F}_{p}$. If $X$ is non-compact but $X^{T}$ is compact, the Poincaré pairing can still be defined on the localized equivariant cohomology ring $H^{*}_{T}(X)_{\mathrm{loc}}$ from (2.8) via localization without introducing compactly supported cohomology, see [AP93, Section 5.2].
Let $X$ be a symplectic manifold for which its three-pointed ($T$-equivariant) genus 00 Gromov–Witten invariants are defined over $\mathbb{Z}$; this includes the class of closed semipositive symplectic manifolds [MS12, Section 6.4] and the non-compact examples we will consider in this article [BMO11], [MO19].
@@
(2.9) $b_{1}\ast_{T}b_{2}=\sum_{A}\sum_{i_{1},\dots,i_{r}\geq 0}\langle b_{1},b_{2},b\rangle_{A}^{i}\ {\lambda}^{i}\ q^{A}\ b^{\vee}\in H^{*}_{T}(X;\Lambda),$
where $\langle b_{1},b_{2},b\rangle_{A}^{i}\ {\lambda}^{i}=\langle b_{1},b_{2},b\rangle_{A}^{i}\ \lambda_{1}^{i_{1}}\cdots\lambda_{r}^{i_{r}}\in H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})$ denotes the (mod $p$ reduction of) three-pointed $T$-equivariant genus 00 Gromov–Witten invariants of degree $A$ with incidence constraints given by Poincaré dual cycles of $b_{1},b_{2},b\in H^{*}_{T}(X;\mathbb{F}_{p})$. The sum is well-defined by Gromov compactness. Extending the product linearly for the Novikov parameters $q^{A}\in\Lambda$, we obtain a product on $H^{*}_{T}(X;\Lambda)$ which is associative and commutative.

### v3

@@ 1.2. Main results
Fix an odd prime $p>2$, and take a cohomology class $b\in H^{*}(X;\mathbb{F}_{p})$. The quantum Steenrod operations [Fuk97] [Wil20] are defined from $\mathbb{Z}/p$-equivariant counts of (perturbed) holomorphic curves $u:\mathbb{P}^{1}\to X$ from $\mathbb{P}^{1}$ with $p$ marked points along the roots of unity, all incident to (Poincaré dual of) $b$. Following [Lee23], we generalize the quantum Steenrod operations to a setting where the target symplectic manifold $(X,\omega)$ admits Hamiltonian torus actions by $\mathbf{T}=(S^{1})^{r}$. For a fixed (equivariant) cohomology class $b\in H^{*}_{\mathbf{T}}(X;\mathbb{F}_{p})$ we define a $\mathbf{T}$-equivariant quantum Steenrod operation (2.17) which is an endomorphism
(1.1) $\Sigma_{b}^{\mathbf{T}}:H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}$
of the underlying module of the equivariant quantum cohomology; here $\Lambda$ is the Novikov ring and $t,\theta$ represent the $\mathbb{Z}/p$-equivariant parameters $H^{*}_{\mathbb{Z}/p}(\mathrm{pt};\mathbb{F}_{p})\cong\mathbb{F}_{p}{[\![t,\theta]\!]}$ of degrees $|t|=2,|\theta|=1$ (since $p>2$, note that $\theta^{2}=0$).
In the presence of a torus $\mathbf{T}$ acting on $X$, there is another set of operations on equivariant quantum cohomology known as the shift operators (2.27) for each choice of a cocharacter $\beta:S^{1}\to\mathbf{T}$,
(1.2) $\mathbb{S}_{\beta}:H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}.$
The key property of the shift operator is that it does not give linear map of $H^{*}_{\mathbf{T}}(\mathrm{pt}){[\![t,\theta]\!]}$-modules, but rather satisfies a “twisted-linear” condition (2.58). Following [Iri17] and [LJ20], we define the shift operator as a composition of a linear map $S_{\beta}$ (the “bare” shift operators) and a reparametrization map accounting for the failure of linearity (and hence the shift), see 2.27.
Our first main result is the following new compatibility relation of the (equivariant) quantum Steenrod operations $\Sigma_{b}^{\mathbf{T}}$ with the shift operators $\mathbb{S}_{\beta}$.
@@
The quantum Steenrod operations $\Sigma_{b}^{\mathbf{T}}$ and the shift operators $\mathbb{S}_{\beta}$ commute. Equivalently, the bare shift operators $S_{\beta}$ commute with $\Sigma_{b}^{\mathbf{T}}$ up to a shift in the equivariant parameters:
(1.3) $\Sigma_{b}^{\mathbf{T}}|_{\lambda\mapsto\lambda-\beta t}\circ S_{\beta}=S_{\beta}\circ\Sigma_{b}^{\mathbf{T}}.$
@@
(1.4) $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}:H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}\to H^{*}_{\mathbf{T}}(X;\Lambda){[\![t,\theta]\!]}.$
@@
Under the hypothesis (3.23) that our analytically defined shift operators agree with the shift operators defined in the algebraic geometry literature, we prove the following.
@@
Consider the action of $\mathbf{T}=T\times S^{1}$ on $T^{*}(G/B)$, as described above. For almost all primes $p$, the $p$-curvature $F_{b}$ and the quantum Steenrod operation $\Sigma_{b}^{\mathbf{T}}$ for $b\in H^{2}(X;\mathbb{F}_{p})$ agree.
Remark 1.3.
The $p$-curvature endomorphism $F_{b}$ is only defined for classes $b\in H^{2}(X;\mathbb{F}_{p})$ of degree $2$, but the quantum Steenrod operation $\Sigma_{b}^{\mathbf{T}}$ can be defined for classes of arbitrary cohomological degree. When the quantum cohomology algebra is generated by classes in $H^{2}$, 1.2 characterizes the quantum Steenrod operation for any class in terms of iterations of the $p$-curvature via the quantum Cartan relation $\Sigma_{b}^{\mathbf{T}}\circ\Sigma_{b^{\prime}}^{\mathbf{T}}=\pm\Sigma_{b\ast b^{\prime}}^{\mathbf{T}}$, where $\ast$ denotes the quantum product. ∎
@@
In particular, the same result with the same proof holds for e.g. generalized Springer resolutions $T^{*}(G/P)$ for parabolic $P\leq G$, Nakajima quiver varieties of type $A$, hypertoric varieties, and Hilbert scheme of points in $\mathbb{C}^{2}$; in general it would hold under the assumptions that (i) $X^{T}$ is discrete and (ii) the quantum multiplication $b\ \ast$ has simple spectrum. ∎
@@ 1.3. Methods
(1.5) $\Sigma_{b}^{\mathbf{T}}|_{\lambda\mapsto\lambda-\beta t}\circ S_{\beta}=S_{\beta}\circ\Sigma_{b}^{\mathbf{T}}$
@@
From 1.1, we deduce 1.2. The key observations are the following (2.18, 4.6) that
@@
To show that such an expansion of $F_{b}$ and $\Sigma_{b}$ as polynomials of $h$ is well-defined, we work in a basis of equivariant cohomology provided by the theory of cohomological stable envelopes of [MO19]. For the case of $X=T^{*}(G/B)$, the stable envelopes are given by certain linear combinations of the classes represented by the conormals of Schubert varieties in $G/B$, see [SZ20]. The main technical input is to show the compactness and transversality of the moduli spaces used to define the quantum Steenrod operations (3.16) and shift operators (3.16), when the input and output incidence constraints are given by the stable envelope classes. (See 3.12 that discusses the recent work of [RZ23] which establishes a relevant maximum principle in more generality).
@@ 1.4. Related work
The main result of [SW22] was covariant constancy, which establishes the compatibility of the quantum Steenrod operation $\Sigma_{b}$ with the quantum differential connection. In light of the above discussion, our main result 1.1 can be understood as a companion theorem to [SW22], as it establishes the compatibility of $\Sigma_{b}$ with the difference connection. There are very few operations which are compatible with both the quantum differential connection and the difference connection from the shift operators, and this constrains the quantum Steenrod operations enough to be equivalent to $p$-curvature.
The $p$-curvature is a fundamental invariant of a connection in positive characteristic, and is an object of classical interest, notably featuring in the Grothendieck–Katz $p$-curvature conjecture [Kat72]. In the recent important work of Etingof–Varchenko [EV24b, EV24a], they develop the theory of periodic pencils of flat connections and prove systematic results about the spectrum of $p$-curvature endomorphisms. Our equivariant quantum connections of conical symplectic resolutions are examples of periodic pencils of flat connections, where periodicity arises from the compatibility with the shift operators. For the main result of this paper, the precise results from the theory of [EV24b, EV24a] can actually be bypassed, but we expect their theory to be a key ingredient for the further study of equivariant quantum connections in positive characteristic.
@@ 1.5. Organization of the paper
In Section 4 we review the specific properties of the equivariant quantum connections of symplectic resolutions and their $p$-curvature. We also briefly introduce the theory of periodic pencils of flat connections of Etingof–Varchenko [EV24b], in particular their structural result about the eigenvalues of the $p$-curvature for such connections.
@@ 1.6. Acknowledgements
We would like thank Paul Seidel for his continued encouragement. We thank Shaoyun Bai, Zihong Chen, Hunter Dinkins, Pavel Etingof, Vasily Krylov, Davesh Maulik, Dan Pomerleano, Ben Webster, Alexander Varchenko, and Nicholas Wilkins for stimulating discussions at various stages. We especially thank Pavel Etingof for explaining his ongoing work with Alexander Varchenko and suggesting its crucial relevance with the current project, leading to the main result. Finally, we thank the anonymous referees for extremely helpful comments and suggestions that improved the exposition.
This work was partially supported by MIT Landis fellowship, by the National Science Foundation through grant DMS-1904997, and by the Simons Foundation through grants 6552299 (the Simons Collaboration on Homological Mirror Symmetry) and 256290 (Simons Investigator).
@@ 2.1.3. Equivariant cohomology
again as the model of $E{\mathbb{Z}/p}$, where $\eta\in{\mathbb{Z}/p}$ acts freely by multiplication of $\zeta=e^{2\pi i/p}$. Denote this action of ${\mathbb{Z}/p}$ by $\sigma$. The space $E{\mathbb{Z}/p}=S^{\infty}$ can be equipped with a ${\mathbb{Z}/p}$-CW complex structure with exactly $p$ many $i$-cells $\Delta^{i},\sigma\Delta^{i},\dots,\sigma^{p-1}\Delta^{i}$ in each dimension $i\geq 0$. The $i$-cell $\Delta^{i}$ has a natural compactification with stratified boundary, whose top strata in the boundary are given by the union $\bigsqcup_{j=0}^{p-1}\sigma^{j}\Delta^{i-1}$ of $(i-1)$-cells for $i$ even and $\sigma\Delta^{i-1}-\Delta^{i-1}$ for $i$ odd. The images of $\Delta^{i}$ under the quotient $E{\mathbb{Z}/p}\to B{\mathbb{Z}/p}$ induces a CW-structure on $B{\mathbb{Z}/p}$ with unique $i$-cell for every $i\geq 0$. For $\mathbb{F}_{p}$-coefficients, the images of (the compactification of) $\Delta^{i}$ in $B{\mathbb{Z}/p}$ become cycles, additively generating $H_{*}^{\mathbb{Z}/p}(\mathrm{pt};\mathbb{F}_{p})=H_{*}(B{\mathbb{Z}/p};\mathbb{F}_{p})$.
@@
Then $t,\theta$ generate $H^{*}_{{\mathbb{Z}/p}}(\mathrm{pt},\mathbb{F}_{p})\cong\mathbb{F}_{p}[\![t,\theta]\!]$. Our notation $\mathbb{F}_{p}[\![t,\theta]\!]$ is for graded (or degreewise) completion of the graded algebra $\mathbb{F}_{p}[t,\theta]$ by the ideal generated by $t,\theta$; in particular $\mathbb{F}_{p}[\![t,\theta]\!]$ is isomorphic as graded algebra to $\mathbb{F}_{p}[t,\theta]$.
Additively, the ${\mathbb{Z}/p}$-equivariant cohomology of a point has one generator in each degree $i\geq 0$, which we will denote by
@@
We again take $S^{\infty}$ with the multiplication action of $S^{1}\leq\mathbb{C}^{*}$ as a model of $ES^{1}$, so that $BS^{1}=\mathbb{CP}^{\infty}$ has one $i$-cell for each $i\in 2\mathbb{Z}_{\geq 0}$. Correspondingly, $H^{*}_{S^{1}}(\mathrm{pt};\mathbb{F}_{p})=\mathbb{F}_{p}[\![\lambda_{1}]\!]$ for a choice of the generator $\lambda_{1}\in H^{2}(\mathbb{CP}^{\infty};\mathbb{F}_{p})$, given by the first Chern class of the universal line bundle $\mathcal{O}_{\mathbb{CP}^{\infty}}(1)$. For a higher-dimensional torus $T=(S^{1})^{r}$, take the product $ET=(S^{\infty})^{r}$ with the product action so that $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})=\mathbb{F}_{p}[\![\lambda_{1},\dots,\lambda_{r}]\!].$ When the torus $T$ is fixed, we will denote this as $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p}):=\mathbb{F}_{p}[\![{\lambda}]\!]$, where the variable $\lambda$ is understood as a shorthand for the tuple of parameters $(\lambda_{1},\dots,\lambda_{r})$.
@@
Note that locally finite cycles in $X$ represent cohomology classes in $H^{*}(X;\mathbb{F}_{p})$ by Poincaré duality. The $T$-invariant locally finite cycles in $X$ give rise to a locally finite cycle in the Borel construction $X\times_{T}ET\to BT$ that is fiberwise (i.e. in $X$) the given locally finite cycle, and correspondingly represent cohomology classes in $H^{*}_{T}(X;\mathbb{F}_{p})$. For the running example of $X=T^{*}(G/B)$, it is classical that its cohomology over $\mathbb{Z}$ is additively generated by locally finite cycles given by the unions of conormals to Schubert subvarieties of $G/B$ [Bri05], see Section 3.
@@ 2.1.4. Equivariant quantum cohomology
For closed $X$, the ordinary cohomology of $X$ admits the nondegenerate Poincaré pairing $(\cdot,\cdot):H^{*}(X;\mathbb{F}_{p})\otimes H^{*}(X;\mathbb{F}_{p})\to\mathbb{F}_{p}$ given by $(b,b^{\prime})=\int_{X}b\cup b^{\prime}$. After fixing a basis of $H^{*}(X)$, for a fixed basis element $b\in H^{*}(X;\mathbb{F}_{p})$, we denote its linear dual under the Poincaré pairing by $b^{\vee}\in H^{*}(X;\mathbb{F}_{p})$. If $X$ admits an action of $T$, the Poincaré pairing can be extended to the equivariant cohomology ring $(\cdot,\cdot)_{T}:H^{*}_{T}(X)\otimes H^{*}_{T}(X)\to H^{*}_{T}(\mathrm{pt})$ using the pushforward under the proper map $X\times_{T}ET\to BT$.
If $X$ is non-compact, the Poincaré pairing becomes a pairing between cohomology and compactly supported cohomology, $(\cdot,\cdot):H^{*}(X;\mathbb{F}_{p})\otimes H^{*}_{c}(X;\mathbb{F}_{p})\to\mathbb{F}_{p}$, see [Bri00, Section 1]. If $X$ is non-compact but $X^{T}$ is compact, the Poincaré pairing can still be defined on the localized equivariant cohomology ring $H^{*}_{T}(X)_{\mathrm{loc}}$ from (2.8) via localization without introducing compactly supported cohomology, see [AP93, 5.2].
We place the following assumption about the topology of the target manifold $X$, for simplicity; it is indeed satisfied for the examples considered in this paper.
Assumption 2.2.
The singular cohomology $H^{*}(X;\mathbb{F}_{p})$ is concentrated in even degrees, and the nontrivial cohomology classes $b\in H^{*}(X;\mathbb{F}_{p})$ can be represented as Poincaré duals of (locally finite) cycles in $X$ given by unions of embedded submanifolds. Moreover, assume that $X^{T}$ is nonempty and compact.
Remark 2.3.
By the Serre spectral sequence attached to the Borel construction $X\times_{T}ET\to BT$, the assumption here implies equivariant formality, namely that $H^{*}_{T}(X)$ is freely generated by the non-equivariant cohomology $H^{*}(X)$ as a $H^{*}_{T}(\mathrm{pt})=H^{*}(BT)$-module. In particular, $H^{*}_{T}(\mathrm{pt})$-linear operations on $H^{*}_{T}(X)$ may be defined on $H^{*}(X)$ and extended linearly. ∎
Let $X$ be a symplectic manifold for which its three-pointed ($T$-equivariant) genus 00 Gromov–Witten invariants are defined over $\mathbb{Z}$; this includes the class of closed semipositive symplectic manifolds [MS12, Section 6.4] and the non-compact examples we will consider in this article [BMO11, MO19].
We choose a basis of $H^{*}(X;\mathbb{F}_{p})$ denoted by $\{b\in H^{*}(X)\}$, which also generate $H^{*}_{T}(X)$ as a $H^{*}_{T}(\mathrm{pt})$-module by 2.2.
@@
(2.9) $b_{1}\ast_{T}b_{2}=\sum_{b}\sum_{A}\sum_{i_{1},\dots,i_{r}\geq 0}\langle b_{1},b_{2},b\rangle_{A}^{i}\ {\lambda}^{i}\ q^{A}\ b^{\vee}\in H^{*}_{T}(X;\Lambda),$
where $\langle b_{1},b_{2},b\rangle_{A}^{i}\ {\lambda}^{i}=\langle b_{1},b_{2},b\rangle_{A}^{i}\ \lambda_{1}^{i_{1}}\cdots\lambda_{r}^{i_{r}}\in H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})$ denotes the (mod $p$ reduction of) three-pointed $T$-equivariant genus 00 Gromov–Witten invariants of degree $A$ with incidence constraints given by Poincaré dual cycles of $b_{1},b_{2},b\in H^{*}_{T}(X;\mathbb{F}_{p})$. The first sum is over the chosen basis elements of $H^{*}(X;\mathbb{F}_{p})$. The total sum is well-defined as an element of $H^{*}_{T}(X;\Lambda)$ by Gromov compactness. Extending the product linearly for the Novikov parameters $q^{A}\in\Lambda$, we obtain a product on $H^{*}_{T}(X;\Lambda)$ which is associative and commutative.